Suppose two coins are tossed. Let B be the random variable representing the number of heads that occur find the probability of getting single head


1
Expert's answer
2021-04-19T05:36:40-0400

Wen the two coins are tossed the outcome is maybe they will both give heads or head tail or tail head or tail tail. (HH,HT, TH, TT). The number of total outcome is 4 therefore the probability of getting a single head will be p(HT, TH)

p(B)= 1/4 + 1/4 =2/4

=1/2